OpenStudy (anonymous):

Write an equation in slope intercept form for the line going through 2,7 and -4,9

OpenStudy (amistre64):

first thing to do is find the slope; subtract one point from the other and stack y/x

OpenStudy (anonymous):

Can you please help

OpenStudy (amistre64):

2,7 - -4,9 ------- 6,-2 ; slope = y/x: -2/6 do you agree?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

then take either point and load up the form: y = mx -mPx+Py

OpenStudy (amistre64):

y = -2/6 x +2(2)/6 + 7

OpenStudy (anonymous):

Okay..... That is slope intercept form right?

OpenStudy (amistre64):

that will simplify into it yes ... i left it with some work for you to do ;)

OpenStudy (anonymous):

How old are you ? ;)

OpenStudy (amistre64):

very old

OpenStudy (anonymous):

How much is that? ;) poop

OpenStudy (amistre64):

im older than poop ... yes lol

OpenStudy (amistre64):

i aint got grandkids .... yet